# New models of nonsingular black hole dark matter from limiting curvature

### v1

@@ Abstract
We consider phenomenological models for nonsingular black hole that satisfy the limiting curvature condition (i.e., that have curvatures that are always sub-Planckian in size) while having a more general dependence on the black hole mass than the most studied examples. These models allow black holes to exist while having regulators that are larger than the horizon scale; it has been shown previously that this can lead to observable consequences in an astrophysical setting, for allowed choices for the regulator scale. Noting that substantial horizon-scale modifications of the metric will affect black hole thermodynamics and Hawking radiation, we study these metrics in the context of primordial black hole dark matter. Considering examples with de Sitter and Minkowski cores, respectively, we study the effect of the regulator in these metrics on the allowed black hole mass ranges (or “bands”), the black hole temperature, specific heat and lifetime, and the bounds on the primordial black hole fraction of the total dark matter density from the observed extragalactic gamma ray background.
@@ I Introduction
Primordial black hole dark matter is one such option that has gained some attention in the recent literature [12, 13, 14, 15]. Black holes are predicted by general relativity and are known to exist from astrophysical observations [16, 17, 18, 19, 20, 21], including the gravitational wave signals of black hole mergers [20] and the direct imaging of black hole shadows [21]. It is not a radical extrapolation to assume that black holes may occur with masses much less than the solar mass scale, and there are known mechanisms for the production of primordial black holes in the early universe [22, 23]. Black hole dark matter does not require the existence of an (unobserved) particle dark sector, or (unobserved) mediator particles that are needed in models that rely on the freeze-out mechanism to obtain the present dark matter relic density. In that sense, it is a scenario that is quite economical. In this paper, we consider a specific class of black holes as possible candidates for primordial black hole dark matter.
@@
The observational implications of nonsingular black holes with astrophysical masses have been considered in the literature [36, 37, 38, 39, 40], as well the possibility that nonsingular black holes may be primordial black hole dark matter [41, 42, 43]. Since the masses of primordial black holes are much smaller than their stellar cousins, and the curvatures near the horizon are correspondingly larger, one might expect that modifications of general relativity will have more important effects in the vicinity of such primordial black holes.
In the present work, we focus on nonsingular black holes that have not been considered previously as candidates for primordial black hole dark matter and that are more general than the most-studied examples. The metrics we consider are guided by the limiting curvature condition [44, 45, 46], which elevates the conjecture that curvature invariants may not exceed maximum values set by the Planck scale to a universal principle. [^1] In fact, limiting curvature has been implemented at the Lagrangian level in general relativity and regular black hole solutions have been found [47]. For concreteness, consider the well-known Hayward metric [48]
@@
where the chosen form of the dimensionless function $f(\hat{\ell})$ defines a given model. This leads to the following finite curvature invariants for $r\ll\ell$,
@@
where $R$ is the Ricci scalar, $C{}_{\mu\nu\rho\sigma}$ denotes the Weyl tensor, and the invariant $S^{2}$ is the square of the tracefree Ricci tensor $S{}_{\mu\nu}=R{}_{\mu\nu}-(1/4)R\,g{}_{\mu\nu}$. The new black hole metrics that are defined via the introduction of the function $f(\hat{\ell})$ are subject only to the requirement that $\ell f(\hat{\ell})^{1/2}$ remains larger that the Planck length for any value of the mass $M$. For example, the choice
@@
has the desired properties. Imagining for the moment that $\epsilon\ll 1$, $f(\hat{\ell})\rightarrow 1+\epsilon\approx 1$ as $M\rightarrow\infty$, while it uniformly decreases to $\epsilon$ as $M\rightarrow 0$. As long as $\sqrt{\epsilon}\,\ell>\ell_{P}$, the limiting curvature condition is satisfied for any $M$ [^2] The $\epsilon$ factor was omitted from the discussion of Ref. [49]; $\epsilon$ may be taken small enough to have a negligible effect on the astrophysical observables considered in Ref. [49] while still satisfying the limiting curvature condition. To see this, we note that despite the explicit $M$ dependence in Eqs. (4) and (5), the maximum values of the curvature invariants are functions of $f(\hat{\ell})\,\ell^{2}$ only:
@@
The maxima are characterized by the length scale $f(\hat{\ell})^{1/2}\,\ell$, up to constants of $\mathcal{O}(1)$, and have only a very mild $M$-dependence via $f(\hat{\ell})$.
The introduction of the function $f(\hat{\ell})$ generalizes a mass dependence that is already present in the quantity $L^{3}$ that appears in the Hayward metric, but is a less conventional choice since it also introduces a mass dependence in the curvature invariants evaluated at $r=0$. In the absence of a rigorous derivation of regular black hole metrics from an underlying theory, it seems reasonable to allow for this possibility, provided that the modifications are consistent with the limiting curvature condition for all $M$. We will also consider a similar modification of a nonsingular black hole metric that has vanishing curvature at the origin, one with phenomenological consequences that are similar to the nonsingular black holes that were discussed in Ref. [49]; in the example we study, the curvature invariants away from the origin satisfy the limiting curvature condition for all $M$. One of the interesting features of the allowed black hole mass bands in the models we consider is that the black holes will Hawking radiate only until their mass hits the edge of a band where its temperature vanishes and the black hole becomes a compact, horizonless object. This will have implications for primordial black hole dark matter as we will discuss later in the present work.
@@ II Nonsingular Black Hole Models
One of the main points of Ref. [49] was that deviations from general relativity can be negligible for masses relevant in precise terrestrial experiments, but may nonetheless be significant for astrophysical black holes, due to the nontrivial mass dependence in $f(\hat{\ell})$. This allows the regulator $\ell$ to be, for example, kilometer-scale and comparable to the horizon scale for astrophysical black holes, leading to observable consequences. For example, the effect on the black hole shadow was found to be up ${\cal O}(5\%)$ relative to the expectation for a Schwarzschild black hole [49].
@@ II.2 Minkowski cores
$F(r)=1-\frac{2GM}{r}\exp\left[-\frac{(2GM\ell^{2})^{1/3}}{r}\,f(\hat{\ell})\right]\,,$ (14)
where $f(\hat{\ell})$ is again given by Eq. (6). Computing the curvature invariants $R$, $C^{2}$, and $S^{2}$ close to the origin $r=0$ we find
@@
These again are only mildly $M$-dependent via the function $f(\hat{\ell})$. Similar to the case of the de Sitter core, a black hole horizon only exists provided
^$\hat{\ell}^{2/3}f(\hat{\ell})\leq\frac{1}{e}\approx 0.3679\,.$ (21)
@@
where again the lower band emerges for $f(\hat{\ell})\not=1$. Note that if one were to replace the quantity $(2GM\ell^{2})^{1/3}f(\hat{\ell})$ in Eq. (14) by a mass-independent parameter with units of length, one would obtain the nonsingular black hole metric that has been studied previously in Refs. [50, 51, 52, 53]; the mass dependence that we have introduced in Eq. (14) not only assures that the limiting curvature condition is valid for all $M$, but also leads to the possibility of large regulator values and new black hole mass ranges, similar to the de Sitter core model defined by Eqs. (1), (2) and (6). See Fig. 2 for a visualization of this band structure. As in our earlier de Sitter core model, the femtometer-sized regulators that are relevant for primordial black hole dark matter do not lead to observable deviations from Newtonian gravity in the laboratory or at larger length scales.
@@ III Thermodynamics and lifetimes
[…]
@@
[…]
@@
Notably, in the case where $\epsilon$ is nonzero, we find that the edges of the black hole bands generically feature vanishing Hawking temperatures with vanishing heat capacity. Moreover, the heat capacity approaches zero from above at the lower end of a given black hole band, and from below at the upper end of a black hole band. This implies that the black holes at the low-mass end of a black hole band can be interpreted as thermodynamically stable end products of Hawking evaporation from larger black holes from within the same band. [^4] The potential role of such remnants as a part of dark matter has recently been addressed in Ref. [55]. This is to be starkly contrasted with the case of Schwarzschild black holes, which always decay, and which exist for all masses. Similarly, the more conventional de Sitter core or Minkowski core black holes discussed in the literature do not feature low-mass bands. Significantly, for $f\not=1$ and $\epsilon\not=0$ the mass range of thermodynamically stable and potentially viable dark matter mass windows is appreciably enlarged.
@@
where $M_{0}$ is the initial black hole mass and ^$\hat{\ell}_{0}\equiv c^{2}\ell/(2GM_{0})$. Equation (30) can easily be integrated numerically. We thereby extract the mass evolution curves depending on the initial mass $M_{0}$ of the black hole; see Fig. 5. In this figure we keep ^$\hat{\ell}_{0}$ fixed, implying that the depicted curves correspond to slightly different magnitudes for the regulator $\ell$. While this choice has little effect in this figure, it is made for consistency with our analysis in Sec. IV: there we restrict our attention to a slice of parameter space with constant $\hat{\ell}$, where black holes over a wider range of masses are expected to have regulator effects at the horizon that are of comparable importance.
@@
Figure 5: Given a choice of initial regulator ^$\hat{\ell}_{0}=\ell/(2GM_{0})$ and black hole model, we plot the evolution of the normalized black hole mass as a function of dimensionless time (measured in the age of the Universe $t_{0}$). The left panel shows the case of de Sitter core black holes with ^$\hat{\ell}_{0}=4$ and $\epsilon=0.001$, and the right panel depicts the case of the Minkowski core black holes with ^$\hat{\ell}_{0}=2$ and $\epsilon=0.001$. For the masses greater than approximately $10^{15}$ g, which are considered throughout the rest of this paper, the black hole mass loss is hence small enough that their masses may be considered constant over the lifetime of the Universe.
@@ IV Dark Matter Fraction
working in units where $\hbar$, $c$ and the Boltzmann constant $k_{B}$ have been set to $1$. The quantity $n_{i}$ is the number of degrees of freedom of the particle in question (for example, $n_{i}=2$ for a massless gauge field with two physical polarizations) and the sign in the denominator is -- ($+$) for bosons (fermions). The greybody factors $\Gamma^{s}_{l,m}(\omega)$, are a function of the particle energy (i.e., $\omega=E_{i}/\hbar$ in conventional units) which captures how the spectrum measured asymptotically far from the black hole horizon differs from that of a blackbody; $l$ and $m$ are the angular momentum labels for the partial wave decomposition of this radiation. The greybody factors may be related to the solutions of the radial Teukolsky equation [56], which determines how the massless perturbations of different spins propagate in the black hole background. We will focus on the photons that are emitted via Hawking radiation. While the computation of the greybody factors is clearly presented in Ref. [41], there are a few typographical errors in some of the equivalent forms of the radial Teukolsky equation that are displayed in that reference; to avoid any confusion, we present the formulae necessary for computing the greybody factors in appendix A using our notation.
Denoting quantities that are measured in the present epoch by the subscript 00, the present number density of photons with energy $E_{\gamma 0}$ may be written
@@
As an example, we show in Fig. 6 the photon flux for a range of de Sitter core black hole masses, corresponding to the choice of $f(\hat{\ell})$ in Eq. (6) with $\epsilon=0.001$. The curve corresponding to each choice of $M$ has $n_{\text{pbh}}(t_{0})$ adjusted so that the flux does not exceed the $1\,\sigma$ error bars on any data point for the observed flux from three observatories (the same considered in Ref. [41]), namely HEAO-1 [58], COMPTEL [59] and EGRET [60]. The analogous plot for the Minkowski core black holes (with $\epsilon=0$) is qualitatively similar, but is not shown. The resulting upper bounds on $n_{\text{pbh}}(t_{0})$ may be used to determine the excluded regions of the $f_{\text{pbh}}$ versus $M$ plane. Characteristic examples are shown in Fig. 7: The left panel compares the excluded fraction for the modified de Sitter core black holes, previously defined, to the $\ell=0$ Schwarzschild limit. The plot assumes ^$\hat{\ell}=4$, with the difference between the dot-dashed and dotted lines indicating how the bound changes as the parameter $\epsilon$ is varied from $0.001$ (dot-dashed line) to zero (dotted line). The right panel compares the excluded fraction for modified Minkowski core black holes, previously defined, to the $\ell=0$ Schwarzschild limit. The bound for the Minkowski core black hole is indicated by the dashed line, assuming ^$\hat{\ell}=2$; the dot-dashed line is again the de Sitter core black hole with ^$\hat{\ell}=4$ and $\epsilon=0.001$, which is provided for comparison. These plots show that a larger fraction of primordial nonsingular black holes are allowed at a given mass compared to the Schwarzschild case, similar to the other nonsingular black holes that were studied in Ref. [41]. The effect of the regulator is less pronounced in the Minkowski core case due to the faster fall off of $(2GM\ell^{2})^{1/3}f(\hat{\ell})$ at large $\ell$ compared to $\ell^{2}f(\hat{\ell})$ in the de Sitter core case.
Figure 7: Excluded values (shaded) of the fraction of dark matter that is made up of primordial black holes. The left panel considers the de Sitter core case with ^$\hat{\ell}=4$; the dot-dashed and dotted lines correspond to the choices $\epsilon=0.001$ and 00, respectively. The right panel shows the excluded regions for the Minkowski core case with ^$\hat{\ell}=2$ (dashed line). The dot-dashed line again represents the de Sitter core ^$\hat{\ell}=4$, $\epsilon=0.001$ line from the left panel, provided for easier comparison.

### v2

@@ Abstract
We consider phenomenological models for nonsingular black holes that satisfy the limiting curvature condition (i.e., that have curvatures that are always sub-Planckian in size) while having a more general dependence on the black hole mass than the most studied examples. These models allow black holes to exist while having regulators that are larger than the horizon scale; it has been shown previously that this can lead to observable consequences in an astrophysical setting, for allowed choices for the regulator scale. Noting that substantial horizon-scale modifications of the metric will affect black hole thermodynamics and Hawking radiation, we study these metrics in the context of primordial black hole dark matter. Considering examples with de Sitter and Minkowski cores, respectively, we study the effect of the regulator in these metrics on the allowed black hole mass ranges (or “bands”), the black hole temperature, specific heat and lifetime, and the bounds on the primordial black hole fraction of the total dark matter density from the observed extragalactic gamma ray background.
@@ I Introduction
Primordial black hole dark matter is one such option that has gained some attention in the recent literature [12, 13, 14, 15, 16]. Black holes are predicted by general relativity and are known to exist from astrophysical observations [17, 18, 19, 20, 21, 22], including the gravitational wave signals of black hole mergers [21] and the direct imaging of black hole shadows [22]. It is not a radical extrapolation to assume that black holes may occur with masses much less than the solar mass scale, and there are known mechanisms for the production of primordial black holes in the early universe [23, 24, 25]. Black hole dark matter does not require the existence of an (unobserved) particle dark sector, or (unobserved) mediator particles that are needed in models that rely on the freeze-out mechanism to obtain the present dark matter relic density. In that sense, it is a scenario that is quite economical. In this paper, we consider a specific class of black holes as possible candidates for primordial black hole dark matter.
@@
The observational implications of nonsingular black holes with astrophysical masses have been considered in the literature [38, 39, 40, 41, 42], as well as the possibility that nonsingular black holes may be primordial black hole dark matter [43, 44, 45, 46] or drivers of cosmic acceleration [47]. Some nonsingular black holes have been developed as an alternative to those with a minimum spherical radius [48, 49], and their quasinormal ringing has also been studied [50]. Since the masses of primordial black holes can be much smaller than the masses of astrophysical black holes, and the curvatures near the horizon are correspondingly larger, one might expect that modifications of general relativity will have more important effects in the vicinity of such primordial black holes.
In the present work, we focus on nonsingular black holes that have not been considered previously as candidates for primordial black hole dark matter and that are more general than the most-studied examples. The metrics we consider are guided by the limiting curvature condition [51, 52, 53, 54], which elevates the conjecture that curvature invariants may not exceed maximum values set by the Planck scale to a universal principle. [^1] Limiting curvature has been implemented at the Lagrangian level in general relativity [55] and nonsingular black hole solutions have been found in $(1+1)$ dimensions [56, 57, 58] and $(3+1)$ dimensions [59]. For concreteness, consider the well-known Hayward metric [60]
@@
where the chosen form of the dimensionless function f(^)f($\hat{\ell}$) defines a given model. This leads to the following finite curvature invariants for $r\ll\ell$,
@@
where $R$ is the Ricci scalar, $C{}_{\mu\nu\rho\sigma}$ denotes the Weyl tensor, and the invariant $S^{2}$ is the square of the tracefree Ricci tensor $S{}_{\mu\nu}=R{}_{\mu\nu}-(1/4)R\,g{}_{\mu\nu}$. The new black hole metrics that are defined via the introduction of the function f(^)f($\hat{\ell}$) are subject only to the requirement that $\ell f(\hat{\ell})^{1/2}$ remains larger that the Planck length for any value of the mass $M$. For example, the choice
@@
has the desired properties. Imagining for the moment that $\epsilon\ll 1$, $f(\hat{\ell})\rightarrow 1+\epsilon\approx 1$ as $M\rightarrow\infty$, while it uniformly decreases to $\epsilon$ as $M\rightarrow 0$. As long as $\sqrt{\epsilon}\,\ell>\ell_{P}$, the limiting curvature condition is satisfied for any MM. [^2] The $\epsilon$ factor was omitted from the discussion of Ref. [61]; $\epsilon$ may be taken small enough to have a negligible effect on the astrophysical observables considered in Ref. [61] while still satisfying the limiting curvature condition. To see this, we note that despite the explicit $M$ dependence in Eqs. (4) and (5), the maximum values of the curvature invariants are functions of $f(\hat{\ell})\,\ell^{2}$ only:
@@
The maxima are characterized by the length scale $f(\hat{\ell})^{1/2}\,\ell$, up to constants of $\mathcal{O}(1)$, and have only a very mild $M$-dependence via f(^)f($\hat{\ell}$).
The introduction of the function f(^)f($\hat{\ell}$) generalizes a mass dependence that is already present in the quantity $L^{3}$ that appears in the Hayward metric, but is a less conventional choice since it also introduces a mass dependence in the curvature invariants evaluated at $r=0$. In the absence of a rigorous derivation of regular black hole metrics from an underlying theory, it seems reasonable to allow for this possibility, provided that the modifications are consistent with the limiting curvature condition for all $M$. We will also consider a similar modification of a nonsingular black hole metric that has vanishing curvature at the origin, one with phenomenological consequences that are similar to the nonsingular black holes that were discussed in Ref. [61]; in the example we study, the curvature invariants away from the origin satisfy the limiting curvature condition for all $M$. One of the interesting features of the allowed black hole mass bands in the models we consider is that the black holes will Hawking radiate only until their mass hits the edge of a band where its temperature vanishes and the black hole becomes a compact, horizonless object. This will have implications for primordial black hole dark matter as we will discuss later in the present work.
@@ II Nonsingular Black Hole Models
In this section, we define the nonsingular black holes of interest, which have either de Sitter or Minkowski cores. Nonsingular de Sitter-core black holes are a generic consequence of the quantum field theory corrections in curved spacetime to the Schwarzschild metric [62, 63]. Mechanisms that generate a positive effective vacuum energy at high curvature arise naturally in many semiclassical and quantum-gravity-motivated models and include quantum phase transitions [64, 65], condensate formation [66], or loop-quantum-gravity-induced corrections [67, 68]. While nonsingular black holes with anti-de Sitter (AdS) cores represent another logical possibility, they have been less frequently studied. AdS cores require sustained negative energy densities and finely tuned matching layers to remain static, which typically demands exotic matter sources not supported by known physics and are therefore considered less plausible [69, 70, 71]. Hence, we do not consider that possibility in the present work. However, we do consider a nonsingular black hole with a Minkowski core, as an example where the limiting curvature condition is satisfied everywhere with vanishing curvature invariants at the origin. One should keep in mind that realistic physical realizations of Minkowski-core black holes may require fine tuning that may not be needed in models where the energy density near the origin is positive.
@@
One of the main points of Ref. [61] was that deviations from general relativity can be negligible for masses relevant in precise terrestrial experiments, but may nonetheless be significant for astrophysical black holes, due to the nontrivial mass dependence in f(^)f($\hat{\ell}$). This allows the regulator $\ell$ to be, for example, kilometer-scale and comparable to the horizon scale for astrophysical black holes, leading to observable consequences. For example, the effect on the black hole shadow was found to be up ${\cal O}(5\%)$ relative to the expectation for a Schwarzschild black hole [61].
@@ II.2 Minkowski cores
where f(^)f($\hat{\ell}$) is again given by Eq. (6). Computing the curvature invariants $R$, $C^{2}$, and $S^{2}$ close to the origin $r=0$ we find
$R$ $=\exp\left[-\frac{(2GM\ell^{2})^{1/3}}{r}\,f(\hat{\ell})\right]\frac{(2GM)^{5/3}f(\hat{\ell})^{2}\ell^{4/3}}{r^{5}}\qquad\text{(exact)}\,$, (15)
@@
These again are only mildly $M$-dependent via the function f(^)f($\hat{\ell}$). Similar to the case of the de Sitter core, a black hole horizon only exists provided
$\hat{\ell}^{2/3}f(\hat{\ell})\leq\frac{1}{e}\approx 0.3679\,.$ (21)
@@
[…]
@@ III Thermodynamics and lifetimes
$r_{H}=\frac{2GM}{3}\left\{1+2\cos\left[\frac{1}{3}\arccos\left(1-\frac{27}{2}\hat{\ell}^{2}f(\hat{\ell})\right)\right]\right\}\quad\text{(de\,Sitter core )}\,$, (25)
@@
$r_{H}=2GM\,\exp\left\{W\left[-\hat{\ell}^{2/3}f(\hat{\ell})\right]\right\}\,,\quad\text{(Minkowski core)}\,$, (26)
@@
Notably, in the case where $\epsilon$ is nonzero, we find that the edges of the black hole bands generically feature vanishing Hawking temperatures with vanishing heat capacity. Moreover, the heat capacity approaches zero from above at the lower end of a given black hole band, and from below at the upper end of a black hole band. This implies that the black holes at the low-mass end of a black hole band can be interpreted as thermodynamically stable end products of Hawking evaporation from larger black holes from within the same band. [^4] The potential role of such remnants as a part of dark matter has been addressed in e.g. Refs. [77, 78]. This is to be starkly contrasted with the case of Schwarzschild black holes, which always decay, and which exist for all masses. Similarly, the more conventional de Sitter core or Minkowski core black holes discussed in the literature do not feature low-mass bands. Significantly, for $f\not=1$ and $\epsilon\not=0$ the mass range of thermodynamically stable and potentially viable dark matter mass windows is appreciably enlarged.
@@
where $M_{0}$ is the initial black hole mass and $\hat{\ell}_{0}\equiv c^{2}\ell/(2GM_{0})$. Equation (30) can easily be integrated numerically. We thereby extract the mass evolution curves depending on the initial mass $M_{0}$ of the black hole; see Fig. 5. In this figure we keep $\hat{\ell}_{0}$ fixed, implying that the depicted curves correspond to slightly different magnitudes for the regulator $\ell$. While this choice has little effect in this figure, it is made for consistency with our analysis in Sec. IV: there we restrict our attention to a slice of parameter space with constant $\hat{\ell}$, where black holes over a wider range of masses are expected to have regulator effects at the horizon that are of comparable importance.
@@
Figure 5: Given a choice of initial regulator $\hat{\ell}_{0}=\ell/(2GM_{0})$ and black hole model, we plot the evolution of the normalized black hole mass as a function of dimensionless time (measured in the age of the Universe $t_{0}$). The left panel shows the case of de Sitter core black holes with $\hat{\ell}_{0}=4$ and $\epsilon=0.001$, and the right panel depicts the case of the Minkowski core black holes with $\hat{\ell}_{0}=2$ and $\epsilon=0.001$. For the masses greater than approximately $10^{15}$ g, which are considered throughout the rest of this paper, the black hole mass loss is hence small enough that their masses may be considered constant over the lifetime of the Universe.
@@ IV Dark Matter Fraction
working in units where $\hbar$, $c$ and the Boltzmann constant $k_{B}$ have been set to $1$. The quantity $n_{i}$ is the number of degrees of freedom of the particle in question (for example, $n_{i}=2$ for a massless gauge field with two physical polarizations) and the sign in the denominator is - (++) for bosons (fermions). The greybody factors $\Gamma^{s}_{l,m}(\omega)$, are a function of the particle energy (i.e., $\omega=E_{i}/\hbar$ in conventional units) which captures how the spectrum measured asymptotically far from the black hole horizon differs from that of a blackbody; $l$ and $m$ are the angular momentum labels for the partial wave decomposition of this radiation. The greybody factors may be related to the solutions of the radial Teukolsky equation [79], which determines how the massless perturbations of different spins propagate in the black hole background. We will focus on the photons that are emitted via Hawking radiation. While the computation of the greybody factors is clearly presented in Ref. [43], there are a few typographical errors in some of the equivalent forms of the radial Teukolsky equation that are displayed in that reference; to avoid any confusion, we present the formulae necessary for computing the greybody factors in appendix A using our notation.
Denoting quantities that are measured in the present epoch by the subscript 0, the present number density of photons with energy $E_{\gamma 0}$ may be written
@@
[…]
Figure 7: Excluded values (shaded) of the fraction of dark matter that is made up of primordial black holes. The left panel considers the de Sitter core case with $\hat{\ell}=4$; the dot-dashed and dotted lines correspond to the choices $\epsilon=0.001$ and 0, respectively. The right panel shows the excluded regions for the Minkowski core case with $\hat{\ell}=2$ (dashed line). The dot-dashed line again represents the de Sitter core $\hat{\ell}=4$, $\epsilon=0.001$ line from the left panel, provided for easier comparison.